IF ALPHA AND BETA ARE THE ROOTS OF THE EQUATION X^2+X+1=0, THEN THE EQUATION WHOSE ROOTS ARE ALPHA^19, BETA^7 IS_______

IF ALPHA AND BETA ARE THE ROOTS OF THE EQUATION X^2+X+1=0,  THEN THE EQUATION WHOSE ROOTS ARE ALPHA^19, BETA^7 IS_______

Grade:11

1 Answers

abhivyakti
6 years ago
x2 +  x+1=0(answer)
the roots of x2 +  x+1=0 are both imaginary ꙍ, ꙍ2.
Alpha19= ꙍ19= ꙍ
Beta7= ꙍ14= ꙍ2
Therefore, the equation whose roots are ꙍ, ꙍ2 is x2+x+1=0